OpenStudy (anonymous):

tiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtired

OpenStudy (bahrom7893):

sqrt(18)/sqrt(3) = sqrt(6*3)/sqrt(3) = sqrt(6)*sqrt(3)/sqrt(3).

OpenStudy (bahrom7893):

you should be able to finish this.

OpenStudy (jhannybean):

\[\large \frac{\sqrt{18}}{\sqrt{3}}\] remember, we can break up 18 into 6 * 3 and \[\large \sqrt{a}* \sqrt{b} = \sqrt{a * b}\] So with that.. \[\large \frac{\sqrt{18}}{\sqrt{3}} = \frac{\sqrt{6}*\sqrt{3}}{\sqrt{3}}\]\[\large \frac{\sqrt{6}*\cancel{\sqrt{3}}}{\cancel{\sqrt{3}}}\] and that leaves you with?